# Learning New Tasks from a Few Examples with Soft-Label Prototypes

### v3

@@ 8 Results and discussion
Our results (Table 1) demonstrate that DeepSLPBERT outperforms BERTfine-tuned in 31/48 tasks, SLPBERT in 43/48 tasks and 1-NN in 41/48 tasks, demonstrating the usefulness of soft-label prototypes and superiority over the “standard” LLM fine-tuning paradigm, as well as the simpler SLP variant. SLPBERT, on the other hand, fares worse than BERTfine-tuned and 1-NN, outperforming the former in only 19/48 tasks and the latter in 17/48 tasks, while exhibiting high standard deviations. Overall, DeepSLP is the best performing method, demonstrating the highest accuracy in 31/48 tasks, while being on-par with the second best model (BERTfine-tuned) on the remaining tasks (15/48 tasks).
SLP vs. DeepSLP
From Table 1, we can see that SLP exhibits high standard deviations, in particular for tasks with a high number of classes ($N>2$). [^5] The only exceptions to this are Disaster ($N=2$), where we observe a high standard deviation across all models; and Emotion ($N=13$) and Political Message ($N=9$), where we observe lower standard deviations. There we only find lines containing 2 centroids (except for one with 3 centroids). This low number of classes per line is therefore responsible for the higher stability and, in turn, lower standard deviations. SLP’s behaviour can be explained by the fact that it can, at times, fail to select the nearest centroid (proof provided in Theorem 12.1), as demonstrated in the example in Figure 5. If we consider the case where $N=2$, then SLP essentially acts as a 1-NN with soft labels trivially at $[1,0]$ and $[0,1]$, with class centroids acting as the nearest neighbour. However, when generalising beyond this setting, the model’s stability is affected. SLP optimises soft labels using the geometric properties of a line and does not consider each (training) data point individually – the soft labels produced by SLP are constants. DeepSLP, on the other hand, learns from training data and produces soft labels as a function of an input point; therefore, it has the ability to output soft labels based on the location of an input (test) point (with the location of the prototypes being fixed).
Figure 5: The centroids for classes A (green), B (pink) and C (red) are located at $(-12,0)$, $(5,0)$ and $(15,0)$ respectively. Unlike most Euclidean distance based algorithms such as 1-NN, the test point marked by $x$ is assigned to B (instead of the nearest class, A) with SLP. The soft labels at $(-15,0)$ and $(15,0)$ are represented in the pie charts on the right.
# SLP DeepSLP
1
$015.6422e-$ 01
$019.7887e-$ 01
$014.3577e-$ 01
$041.3500e-$ 04
$169.7973e-$ 16
$022.0995e-$ 02
2
$135.3090e-$ 13
$022.5240e-$ 02
$014.3212e-$ 01
$051.0000e-$ 05
$015.6787e-$ 01
$019.7475e-$ 01
Table 2: Soft labels derived from SLP and DeepSLP. # denotes the index of the soft-label prototype lying on the line.
In Table 2, we further present the probability distribution of two soft-label prototypes for the category Airline using SLP and DeepSLP with the 16-shot setting. It can be seen that the distribution of soft labels derived from SLP tends to be more skewed, with the probability distribution lying heavily on 2/3 classes, which is in contrast to DeepSLP that seems to be more uniform. This is a pattern that we generally observe across all tasks. This can further explain the lower performance of SLP, as this behaviour essentially tends to dismiss an entire class. DeepSLP, on the other hand, leads to better generalisation and a more stable performance in our highly limited data setting.
Number of shots
For DeepSLP, we observe a general trend where performance increases as the number of shots increases from 4 to 16. On the other hand, for SLP, we do not see as clear a picture, indicating that more data does not necessarily improve performance (e.g., there is a decrease in performance for SLP in $7/16$ test tasks as we move from 8 to 16 shots).
(Deep)SLP training time and model parameters
Training time reduces to three operations: encoding the data, generating lines from class centroids, and constructing soft-label prototypes at the ends of the lines. Generally, we observe that (a) all lines are constructed under a second, (b) soft labels via optimisation of linear constraints using the MOSEK solver are generated in less than a second, and (c) soft labels learned by minimising training loss require less than 10 seconds per line (tasks with $N=2$ or $N=3$ use only one line, except for emotion which has $N=13$ and takes less than a minute). Furthermore, the number of parameters are substantially smaller compared to those in fine-tuning an LLM – we only need to fine-tune $1500-10000$ parameters (based on the number of classes) for the linear layer of each soft-label prototype for DeepSLP, compared to $>10^{8}$ parameters for finetuning BERT. DeepSLP does not require a GPU.
9 Further analysis
9.1 DeepSLP architecture analysis
[…]
9.2 Using a more powerful encoder
While our focus is on extreme few-shot learning utilising limited amounts of training data, in this section, we seek to analyse model behaviour when auxiliary labelled data is available. Specifically, we utilise another encoder, MT-DNN Liu et al. (2019), a BERT model that has been fine-tuned on the GLUE Wang et al. (2018) dataset (similarly to Bansal et al. (2019)), and which we further fine-tune on our data, similarly to our fine-tuned BERT baseline. In Table 4, we present the performance of MT-DNN (fine-tuned) and a DeepSLP variant that uses the MT-DNN encoder (fixed). We can see that the performance of the two is closer in this case (compared to when using the BERT encoder in Table 1) – DeepSLP outperforms MT-DNN in
@@ 9.2 Using a more powerful encoder
Category (Classes) Shot BERT DeepSLPBERT
Rating DVD (3) 4 36.142 ± 8.17 39.566 ± 5.086
8 37.855 ± 5.208 38.788 ± 4.449
16 41.465 ± 4.597 40.53 ± 4.375
Rating Electronics (3) 4 37.815 ± 9.3 39.977 ± 5.959
8 40.901 ± 3.97 41.926 ± 3.985
16 45.289 ± 2.832 44.917 ± 3.164
Political Bias (2) 4 52.466 ± 2.776 53.251 ± 4.042
8 57.819 ± 4.709 58.209 ± 5.198
16 61.025 ± 2.879 61.479 ± 2.974
Rating Kitchen (3) 4 38.954 ± 12.067 39.624 ± 6.787
8 40.511 ± 8.019 41.081 ± 6.777
16 44.923 ± 3.536 45.801 ± 4.562
Political Audience (2) 4 51.069 ± 1.259 51.741 ± 2.827
8 54.117 ± 2.824 54.506 ± 3.274
16 56.827 ± 3.438 56.956 ± 3.045
Sentiment Kitchen (2) 4 58.697 ± 4.943 60.76 ± 4.426
8 63.473 ± 3.207 65.733 ± 3.198
16 69.14 ± 2.413 69.18 ± 2.589
Disaster (2) 4 54.005 ± 10.75 54.252 ± 9.843
8 61.644 ± 6.371 61.3 ± 7.961
16 68.378 ± 2.686 69.28 ± 2.358
Airline (3) 4 44.714 ± 14.082 50.987 ± 4.936
8 53.887 ± 7.074 55.209 ± 6.049
16 59.967 ± 3.842 60.247 ± 4.577
Emotion (13) 4 10.068 ± 2.714 9.706 ± 1.108
8 11.876 ± 0.743 8.041 ± 2.797
16 12.124 ± 1.945 10.919 ± 1.615
Rating Books (3) 4 40.832 ± 8.488 42.116 ± 4.725
8 41.842 ± 5.425 42.156 ± 4.608
16 46.494 ± 2.465 46.513 ± 3.036
Political Message (9) 4 15.311 ± 2.265 14.421 ± 1.095
8 18.244 ± 1.636 16.919 ± 1.756
16 18.971 ± 2.206 18.319 ± 1.74
Scitail (2) 4 51.474 ± 6.275 54.101 ± 3.759
8 55.173 ± 6.254 56.341 ± 5.786
16 59.887 ± 3.667 59.692 ± 4.227
Sentiment DVD (2) 4 54.457 ± 2.97 55.003 ± 2.936
8 56.54 ± 4.009 57.527 ± 3.562
16 60.593 ± 2.73 60.76 ± 2.944
Sentiment Books (2) 4 57.217 ± 3.822 58.67 ± 4.753
8 63.87 ± 3.982 64.78 ± 2.615
16 67.137 ± 3.107 67.453 ± 3.085
Restaurant (8) 4 25.63 ± 3.834 47.634 ± 5.237
8 37.475 ± 2.683 55.912 ± 4.494
16 44.132 ± 3.33 61.716 ± 2.208
CoNLL (4) 4 47.97 ± 5.926 52.724 ± 5.84
8 62.647 ± 4.582 60.374 ± 3.731
16 70.314 ± 2.23 67.496 ± 4.551
Table 3: Contrasting the performance of DeepSLP to a counterpart (BERT) that only tests the architecture of DeepSLP.
23/48 tasks, but the differences are small ($<1\%$ in 18/48 tasks, with similar standard deviation). Again, this highlights DeepSLP’s effectiveness in data lean settings. Furthermore, DeepSLP does not require a GPU while only trains a fraction of the parameters used in MT-DNN finetuning.
Category (Classes) Shot MT-DNN DeepSLPMT-DNN
Rating DVD (3) 4 51.26 ± 8.3 57.87 ± 7.3
8 55.39 ± 10.07 53.96 ± 5.68
16 60.02 ± 1.8 56.2 ± 6.43
Rating Electronics (3) 4 59.82 ± 4.24 57.63 ± 7.91
8 63.37 ± 0.78 60.51 ± 2.91
16 62.85 ± 0.62 61.47 ± 2
Political Bias (2) 4 62.46 ± 4.09 61.25 ± 9.25
8 66.37 ± 0.36 63.83 ± 4.53
16 66.10 ± 3.79 65.63 ± 2.03
Rating Kitchen (3) 4 54.80 ± 13.02 56.81 ± 13.42
8 57.08 ± 11.95 57.47 ± 9.99
16 63.95 ± 4.34 62.02 ± 4.83
Political Audience (2) 4 51.03 ± 7.22 52.87 ± 7.28
8 56.74 ± 5.30 56.78 ± 3.72
16 58.70 ± 2.15 58.78 ± 2.83
Sentiment Kitchen (2) 4 82.50 ± 3.49 86.65 ± 0.54
8 86.67 ± 0.60 86.09 ± 1.26
16 86.63 ± 0.50 86.87 ± 0.26
Disaster (2) 4 50.29 ± 3.60 51.28 ± 3.44
8 51.61 ± 5.07 56.67 ± 5.29
16 56.08 ± 4.77 53.23 ± 1.94
Airline (3) 4 70.41 ± 2.45 63.65 ± 11.7
8 70.22 ± 2.47 70.8 ± 2.69
16 71.04 ± 0.64 72.51 ± 1.14
Emotion (13) 4 12.88 ± 0.73 10.39 ± 2.71
8 15.39 ± 1.37 11.17 ± 2.16
16 17.45 ± 0.56 11.24 ± 1.99
Rating Books (3) 4 65.52 ± 8.96 62.25 ± 6.79
8 69.28 ± 0.27 63.27 ± 3.16
16 68.94 ± 0.78 68.02 ± 1.8
Political Message (9) 4 19.78 ± 1.10 18.38 ± 2.13
8 21.56 ± 0.65 21.16 ± 1.72
16 23.63 ± 2.47 23.02 ± 1.11
Scitail (2) 4 56.11 ± 3.52 63.71 ± 6.85
8 69.36 ± 3.41 73.33 ± 5.63
16 74.18 ± 3.38 73.81 ± 3.66
Sentiment DVD (2) 4 84.30 ± 0.96 84.57 ± 1.05
8 84.90 ± 0.75 84.18 ± 1.41
16 84.73 ± 0.52 85.09 ± 0.28
Sentiment Books (2) 4 85.53 ± 0.99 86.23 ± 0.4
8 85.43 ± 0.34 85.6 ± 1.23
16 85.73 ± 0.71 86.18 ± 0.33
Restaurant (8) 4 28.37 ± 2.65 34.28 ± 5.23
8 46.65 ± 3.80 44.15 ± 5.43
16 58.66 ± 1.24 47.09 ± 3.71
CoNLL (4) 4 35.97 ± 2.38 47.33 ± 6.2
8 49.04 ± 0.63 59.13 ± 5.14
16 64.55 ± 1.14 61.96 ± 5
Table 4: Contrasting the performance of DeepSLP to MT-DNN.
10 Ensemble properties of DeepSLP
In this study, we compare and contrast DeepSLP to ensembles and draw similarities between the two, shedding further light into the effectiveness of our approach. Each prediction decision by DeepSLP is the result of two soft-label prototypes – those that lie on each end of the line nearest to a test point $x$. An analogy can then be drawn between the prototypes used at prediction time and those individual (albeit independent) models that are utilised by an ensemble when producing the final classification. While DeepSLP prototypes are not independent but are rather trained jointly (and share the same encoder), in what follows, we demonstrate that they display several properties of ensemble methods, while being computationally efficient and utilising a small number of parameters. For the analyses below, we consider the tasks Airline and Disaster using an 8-shot setting and evaluate on the test data for each.
10.1 Individual vs joint prediction

### v4

@@ A.7 Results
Our results (Table 5) demonstrate that DeepSLPBERT outperforms BERTfine-tuned in $31/48$ tasks, constraintSLPBERT in $43/48$ tasks and LORABERT in $45/48$ tasks, demonstrating the usefulness of soft-label prototypes and superiority over the “standard” LLM fine-tuning paradigm, as well as the simpler constraintSLP variant. constraintSLPBERT, on the other hand, fares worse than BERTfine-tuned and LORA, outperforming the former in only $19/48$ tasks and the latter in $25/48$ tasks, while exhibiting high standard deviations which can be explained by Theorem A.2, as constraintSLP can behave erratically and not select the closest point to the class centroid. Overall, DeepSLP is the best performing method, demonstrating the highest accuracy in $31/48$ tasks, while being on-par with the second best model (BERTfine-tuned) on the remaining tasks ($15/48$ tasks). Fine-tuned BERT is, overall, the next best model with $13/48$ tasks while constraintSLP achieves the best performance amongst all methods in only $1/48$ tasks. ProtoNet’s comparatively lower performance can be explained by the fact that meta-learning approaches tend to require a large number of diverse and structured meta-training tasks for effective learning — thus not making them readily suited for (extreme) few-shot learning settings.
MetaSLP
In Table 6, MetaSLPREPTILE outperforms all baselines achieving the highest performance in $33/48$ tasks. LEOPARD is the next best model with the highest performance in $11/48$ tasks. Interestingly, MetaSLPFOMAML does not fare as well as MetaSLPREPTILE and achieves the highest performance in only $1/48$ tasks while outperforms LEOPARD in only $6/48$ tasks. MetaSLPFOMAML nevertheless outperforms MetaSLPREPTILE and Reptile in natural language inference tasks – demonstrating the usefulness of learnable inner-loop learning rates across multiple task distributions while meta-training.
@@
Category (Classes) Shot LORABERT ProtoNet constraintSLPBERT BERTfine-tuned* DeepSLPBERT
Political Bias (2) 4 52.75 ± 4.33 51.15 ± 2.454 53.447 ± 3.281 54.57 ± 5.02 53.251 ± 4.042
8 53.66 ± 4.25 56.568 ± 4.228 55.824 ± 3.725 56.15 ± 3.75 58.209 ± 5.198
16 59.21 ± 2.27 59.183 ± 4.706 58.277 ± 4.128 60.96 ± 4.25 61.479 ± 2.974
Emotion (13) 4 7.56 ± 2.93 8.953 ± 2.052 8.662 ± 6.213 09.20 ± 3.22 9.076 ± 1.108
8 9.02 ± 2.36 10.857 ± 3.436 8.16 ± 3.266 08.21 ± 2.12 8.041 ± 2.797
16 10.29 ± 1.67 11.479 ± 2.96 8.115 ± 3.66 13.43 ± 2.51 10.919 ± 1.615
Sentiment Books (2) 4 51.27 ± 2.75 55.53 ± 4.097 59.89 ± 5.385 54.81 ± 3.75 58.67 ± 4.753
8 58.16 ± 3.3 58.97 ± 4.909 64.34 ± 2.565 53.54 ± 5.17 64.78 ± 2.615
16 59.16 ± 2.59 65.5 ± 7.026 66.36 ± 2.183 65.56 ± 4.12 67.453 ± 3.085
Rating DVD (3) 4 31.65 ± 4.91 37.665 ± 7.184 32.298 ± 16.263 32.22 ± 08.72 39.566 ± 5.086
8 37.69 ± 3.16 37.008 ± 5.118 32.644 ± 16.016 36.35 ± 12.50 38.788 ± 4.449
16 38.63 ± 5.52 39.123 ± 6.004 35.587 ± 17.445 42.79 ± 10.18 40.53 ± 4.375
Rating Electronics (3) 4 31.66 ± 2.94 33.696 ± 5.55 35.188 ± 16.211 39.27 ± 10.15 39.977 ± 5.959
8 38.72 ± 5.95 37.297 ± 5.938 29.624 ± 12.876 28.74 ± 08.22 41.926 ± 3.985
16 39.15 ± 6.6 43.825 ± 5.946 29.836 ± 12.753 45.48 ± 06.13 44.917 ± 3.164
Rating Kitchen (3) 4 36.63 ± 4.68 35.914 ± 6.678 28.253 ± 15.907 34.76 ± 11.20 39.624 ± 6.787
8 39.69 ± 6.22 38.46 ± 11.124 24.397 ± 11.961 34.49 ± 08.72 41.081 ± 6.777
16 38.17 ± 7.14 46.546 ± 8.394 31.926 ± 18.29 47.94 ± 08.28 45.801 ± 4.562
Political Audience (2) 4 49.75 ± 1.03 50.976 ± 1.84 51.305 ± 2.68 51.02 ± 1.72 51.741 ± 2.827
8 54.05 ± 2.54 52.022 ± 3.964 53.104 ± 3.669 52.80 ± 2.72 54.506 ± 3.274
16 55.39 ± 3.66 54.024 ± 3.071 53.888 ± 3.305 58.45 ± 4.98 56.956 ± 3.045
Sentiment Kitchen (2) 4 53.02 ± 1.54 55.24 ± 3.427 61.96 ± 4.594 56.93 ± 7.10 60.76 ± 4.426
8 55.54 ± 3.47 62.28 ± 5.103 64.83 ± 3.983 57.13 ± 6.60 65.733 ± 3.198
16 58.59 ± 4.83 66.9 ± 5.441 68.21 ± 3.298 68.88 ± 3.39 69.18 ± 2.589
Disaster (2) 4 56.02 ± 6.35 51.474 ± 8.848 52.77 ± 10.803 55.73 ± 10.29 54.252 ± 9.843
8 57.46 ± 6.9 60.661 ± 4.991 56.888 ± 11.139 56.31 ± 09.57 61.3 ± 7.961
16 65.79 ± 2.03 63.893 ± 6.62 65.907 ± 3.691 64.52 ± 08.93 69.28 ± 2.358
Airline (3) 4 24.36 ± 5.42 44.167 ± 10.752 36.243 ± 22.607 42.76 ± 13.50 50.987 ± 4.936
8 52.31 ± 7.89 50.148 ± 13.429 44.972 ± 22.584 38.00 ± 17.06 55.209 ± 6.049
16 54.1 ± 8.57 54.8 ± 10.49 29.238 ± 17.494 58.01 ± 08.23 60.247 ± 4.577
Rating Books (3) 4 34.69 ± 2.12 37.715 ± 5.801 25.562 ± 15.207 39.42 ± 07.22 42.116 ± 4.725
8 39.36 ± 6.33 38.518 ± 5.327 34.026 ± 14.123 39.55 ± 10.01 42.156 ± 4.608
16 41.23 ± 5.32 44.694 ± 7.797 32.509 ± 16.132 43.08 ± 11.78 46.513 ± 3.036
Political Message (9) 4 12.16 ± 1.46 13.888 ± 2.076 12.438 ± 1.799 15.64 ± 2.73 14.421 ± 1.095
8 15.71 ± 2.04 16.155 ± 2.316 15.08 ± 2.925 13.38 ± 1.74 16.919 ± 1.756
16 15.53 ± 2.55 18.324 ± 2.011 13.121 ± 3.294 20.67 ± 3.89 18.319 ± 1.74
Sentiment DVD (2) 4 50.77 ± 0.78 51.06 ± 3.302 56.06 ± 2.408 54.98 ± 3.96 55.003 ± 2.936
8 52.24 ± 1.54 55.19 ± 3.298 56.98 ± 3.299 55.63 ± 4.34 57.527 ± 3.562
16 52.6 ± 2.09 59.45 ± 3.84 58.95 ± 2.813 58.69 ± 6.08 60.76 ± 2.944
Scitail (2) 4 43.36 ± 4.74 50.227 ± 5.69 52.296 ± 4.366 58.53 ± 09.74 54.101 ± 3.759
8 54.29 ± 5.25 54.196 ± 6.678 55.964 ± 5.705 57.93 ± 10.70 56.341 ± 5.786
16 52.68 ± 3.0 57.744 ± 5.696 59.675 ± 4.033 65.66 ± 06.82 59.692 ± 4.227
Restaurant (8) 4 10.56 ± 1.36 18.161 ± 2.822 24.932 ± 17.102 49.37 ± 4.28 47.634 ± 5.237
8 20.92 ± 2.4 32.146 ± 5.785 29.787 ± 9.573 49.38 ± 7.76 55.912 ± 4.494
16 29.37 ± 4.05 40.435 ± 3.348 29.154 ± 13.537 69.24 ± 3.68 61.716 ± 2.208
CoNLL (4) 4 21.48 ± 2.71 35.438 ± 7.324 27.02 ± 7.346 50.44 ± 08.57 52.724 ± 5.84
8 29.84 ± 3.28 44.259 ± 4.886 31.296 ± 17.487 50.06 ± 11.30 60.374 ± 3.731
16 37.18 ± 3.32 52.116 ± 5.354 22.923 ± 7.933 74.47 ± 03.10 67.496 ± 4.551
Table 5: Classification performance (accuracy) of our methods (constraintSLP and DeepSLP) and baselines in the low-resource setting. Entries in grey indicate the best model out of all; * refers to the baseline as reported in Bansal et al. (2020a). Subscripts for constraintSLP and DeepSLP refer to the (non-fine-tuned) encoder used. Each set of results is separated by a double line. The first set of results contains intent classification tasks, the second set has a natural language inference task and the last set contains entity typing tasks.
Category (Classes) Shot ProtoNet LEOPARD* Reptile MetaSLPREPTILE MetaSLPFOMAML
Political Bias (2) 4 56.33 ± 4.37 60.49 ± 6.66 58.82 ± 4.31 60.96 ± 6.13 55.06 ± 5.9
8 58.87 ± 3.79 61.74 ± 6.73 59.43 ± 3.79 63.65 ± 4.57 58.97 ± 5.5
16 57.01 ± 4.44 65.08 ± 2.14 62.21 ± 0.72 66.05 ± 1.57 63.63 ± 4.74
Emotion (13) 4 09.18 ± 3.14 11.71 ± 2.16 11.65 ± 3.21 11.94 ± 1.95 11.03 ± 2.98
8 11.18 ± 2.95 12.90 ± 1.63 10.56 ± 2.85 13.42 ± 1.46 12.38 ± 2.69
16 12.32 ± 3.73 13.38 ± 2.20 11.62 ± 3.11 14.03 ± 2.35 12.32 ± 1.76
Sentiment Books (2) 4 73.15 ± 5.85 82.54 ± 1.33 76.95 ± 1.03 83.22 ± 0.95 74.51 ± 5.25
8 75.46 ± 6.87 83.03 ± 1.28 77.49 ± 1.08 83.8 ± 0.8 79.25 ± 1.97
16 77.26 ± 3.27 83.33 ± 0.79 77.88 ± 0.56 83.8 ± 1.59 78.41 ± 1.08
Rating DVD (3) 4 47.73 ± 6.20 49.76 ± 9.80 45.91 ± 9.85 45.2 ± 8.91 39.64 ± 5.17
8 47.11 ± 4.00 53.28 ± 4.66 47.23 ± 9.22 58.38 ± 2.9 52.35 ± 5.27
16 48.39 ± 3.74 53.52 ± 4.77 48.49 ± 8.88 57.41 ± 4.71 60.4 ± 3.71
Rating Electronics (3) 4 37.40 ± 3.72 51.71 ± 7.20 44.47 ± 8.25 45.34 ± 7.22 39.53 ± 5.76
8 43.64 ± 7.31 54.78 ± 6.48 49.1 ± 6.81 55.10 ± 5.12 47.83 ± 5.94
16 44.83 ± 5.96 58.69 ± 2.41 50.68 ± 6.8 59.47 ± 2.29 56.53 ± 4.36
Rating Kitchen (3) 4 44.72 ± 9.13 50.21 ± 09.63 45.38 ± 10.96 45.20 ± 8.78 39.11 ± 7.16
8 46.03 ± 8.57 53.72 ± 10.31 46.71 ± 9.84 54.53 ± 9.9 50.19 ± 8.36
16 49.85 ± 9.31 57.00 ± 08.69 52.87 ± 9.52 58.94 ± 7.58 57.63 ± 8.37
Political Audience (2) 4 51.47 ± 3.68 52.60 ± 3.51 52.45 ± 4.26 54.1 ± 3.66 52.03 ± 2.73
8 51.83 ± 3.77 54.31 ± 3.95 52.87 ± 4.31 56.01 ± 3.65 52.06 ± 2.27
16 53.53 ± 3.25 57.71 ± 3.52 55.6 ± 1.85 58.57 ± 2.04 54.33 ± 3.14
Sentiment Kitchen (2) 4 62.71 ± 9.53 78.35 ± 18.36 69.81 ± 14.58 81.96 ± 3.73 72.73 ± 7.97
8 70.19 ± 6.42 84.88 ± 1.12 75.76 ± 1.13 83.33 ± 1.99 76.86 ± 4.46
16 71.83 ± 5.94 85.27 ± 1.31 76.41 ± 0.66 84.33 ± 1.81 80.78 ± 4.38
Disaster (2) 4 50.87 ± 1.12 51.45 ± 4.25 49.76 ± 4.73 55.03 ± 8.73 52.62 ± 2.71
8 51.30 ± 2.30 55.96 ± 3.58 52.17 ± 5.17 57.77 ± 6.40 55.04 ± 5.79
16 52.76 ± 2.92 61.32 ± 2.83 55.37 ± 4.53 65.18 ± 4.41 62.27 ± 4.42
Airline (3) 4 40.27 ± 8.19 54.95 ± 11.81 57.11 ± 14.16 57.39 ± 7.83 51.62 ± 10.53
8 51.16 ± 7.60 61.44 ± 03.90 64.37 ± 3.49 65.67 ± 4.82 57.47 ± 9.37
16 48.73 ± 6.79 62.15 ± 05.56 66.31 ± 2.55 69.48 ± 2.06 65.02 ± 5.16
Rating Books (3) 4 48.44 ± 7.43 54.92 ± 6.18 56.57 ± 8.17 55.79 ± 5.61 54.4 ± 5.83
8 52.13 ± 4.79 59.16 ± 4.13 57.33 ± 7.63 65.74 ± 5.58 57.17 ± 6.77
16 57.28 ± 4.57 61.02 ± 4.19 63.26 ± 3.59 67.87 ± 3.45 $6$ 3.93
Political Message (9) 4 14.22 ± 1.25 15.69 ± 1.57 14.58 ± 1.78 18.84 ± 1.82 14.96 ± 1.94
8 15.67 ± 1.96 18.02 ± 2.32 15.13 ± 2.16 20.09 ± 2.71 16.09 ± 2.6
16 16.49 ± 1.96 18.07 ± 2.41 16.38 ± 2.15 23.22 ± 1.17 16.62 ± 2.19
Sentiment DVD (2) 4 74.38 ± 2.44 80.32 ± 1.02 72.03 ± 11.61 80.97 ± 1.21 73.08 ± 7.56
8 75.19 ± 2.56 80.85 ± 1.23 75.79 ± 1.62 81.85 ± 1.79 76.55 ± 2.9
16 75.26 ± 1.07 81.25 ± 1.41 76.69 ± 0.8 83.48 ± 1.01 78.19 ± 1.32
Scitail (2) 4 76.27 ± 4.26 69.50 ± 9.56 59.13 ± 10.58 53.48 ± 5.59 61.55 ± 9.11
8 78.27 ± 0.98 75.00 ± 2.42 62.63 ± 10.85 60.79 ± 4.6 68.03 ± 4.54
16 78.59 ± 0.48 77.03 ± 1.82 68.03 ± 1.57 61.67 ± 3.61 68.5 ± 3.7
Restaurant (8) 4 17.36 ± 2.75 49.84 ± 3.31 13.37 ± 2.25 27.00 ± 2.61 20.31 ± 2.97
8 18.70 ± 2.38 62.99 ± 3.28 16.83 ± 3.42 35.66 ± 2.39 27.74 ± 2.29
16 16.41 ± 1.87 70.44 ± 2.89 16.0 ± 3.44 37.20 ± 2.68 28.57 ± 2.41
CoNLL (4) 4 32.23 ± 5.10 54.16 ± 6.32 31.31 ± 5.32 40.79 ± 3.40 36.07 ± 3.25
8 34.49 ± 5.15 67.38 ± 4.33 33.17 ± 5.1 41.25 ± 5.21 40.5 ± 2.16
16 33.75 ± 6.05 76.37 ± 3.08 34.04 ± 3.59 45.96 ± 4.75 43.67 ± 6.92
Table 6: Classification performance (accuracy) of MetaSLP and baselines in the high-resource setting. Entries in green indicate the best model out of all; * refers to the baseline as reported in Bansal et al. (2020a). Each set of results is separated by a double line. The first set of results contains intent classification tasks, the second set has a natural language inference task and the last set contains entity typing tasks.
A.8 Ensemble properties of DeepSLP
In this study, we compare and contrast DeepSLP to ensembles and draw similarities between the two, shedding further light into the effectiveness of our approach. Each prediction decision by DeepSLP is the result of two soft-label prototypes – those that lie on each end of the line nearest to a test point $x$. An analogy can then be drawn between the prototypes used at prediction time and those individual (albeit independent) models that are utilised by an ensemble when producing the final classification.
While DeepSLP prototypes are not independent but are rather trained jointly (and share the same encoder), in what follows, we demonstrate that they display several properties of ensemble methods, while being computationally efficient and utilising a small number of parameters. For the analyses below, we consider the tasks Airline and Disaster using an 8-shot setting and evaluate on the test data for each. However, we find the below properties to generalise across all tasks.
A.8.1 Individual vs joint prediction